Shift report 20011013 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Oct 13 2001 - 08:14:07 EDT

  • Next message: Apache: "Shift report 20011013 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Oct 13 08:14:06 2001, Shift: 00:00-08:00
    Supervisor: J.I.Jordre, with: 
    Log book#9 pages 145 - 146
    
    Shift Summary:
    ------------------------------------------------------------
    1. A poor night for the people in MCR. 4 attempted stores between 1 and 3:30,
       but the beam was lost when ramping everytime. Then one quenc link 
       interlock, followed by several attemted stores with only 6 bunches.
       The beam intesity as function of time looks like a bad drawing of the Alps.
       We should send flowers to MCR.
    
    2. DVM, TPC pedestals, DVM and pedestals, runs 5304, 5305, 5306 and 5307 
       respectively.
    
    3. None.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Oct 13 2001 - 08:15:29 EDT